You are considering a project that promises you cash flows of 554 USD each year for 3 years. Based on the riskiness of the project, you require a 12 percent return. What is the maximum you should be willing to pay?

You are considering a project that promises you cash flows of 508 USD each year for 5 years. Based on the riskiness of the project, you require a 12 percent return. The cost to buy into the project is 3,459. What is the project NPV?
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=554[1-(1.12)^-3]/0.12

=554*2.401831268

=USD 1330.61(Approx).

b.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=508[1-(1.12)^-5]/0.12

=508*3.604776202

=1831.23

NPV=Present value of inflows-Present value of outflows

=1831.23-3459

(USD 1627.77)(Negative)(Approx).


answered by: ANURANJAN SARSAM